# Constrained and Ordered Level Planarity Parameterized by the Number of Levels

### v1

@@ 2 An XP / XNLP Algorithm for Ordered Level Planarity
Our table $T$ has $2^{\binom{h}{2}}\prod_{j\in[h]}(2\lambda_{j}+1)$ entries. In view of 1, we can compute the value of a table entry $T[s,U]$ in polynomial time (disregarding the time spent for the recursive calls) by simply constructing the tuples $(s_{1}^{\prime},U_{1}^{\prime}),\dots,(s_{h}^{\prime},U_{h}^{\prime})$ and then setting $T[s,U]=\bigvee_{j\in[h]}T[s_{j}^{\prime},U_{j}^{\prime}]$ (only taking into account the defined quantities). Hence, by employing memoization and the usual back-linked strategy, we obtain the claimed deterministic algorithm. For the nondeterministic algorithm, we perform $2n$ steps. In each step, we nondeterministically guess the next separation $s$ and its set of prescribed edges $U$ and check whether for some $j\in[h]$, the previous separation $s^{\prime}$ is equal to $s_{j}^{\prime}$ and the previous set of prescribed edges $U^{\prime}$ is equal to $U_{j}^{\prime}$ (starting with $s^{\prime}=(0,0,\dots,0)$ and $U^{\prime}=\emptyset$). At each point in time, we only need to keep two separations, two edge sets, and $\mathcal{O}(1)$ pointers in memory, which can be done using $\mathcal{O}(h^{2}+h\log I)$ space.
As a corollary of Lemmas 2.1, 2.2 and 2.4, we obtain the algorithmic statements in Theorem 1.1.
@@ 3 XNLP-Hardness of Ordered Level Planarity
Our construction of $\mathcal{G}$ is heavily based on two gadgets that we call plugs and sockets (a very basic version of these gadgets was already used earlier, in an NP-hardness proof for CLP [7]; here we introduce generalized versions). We define both in terms of the list of levels their vertices occupy. Their $\prec_{i}$ orderings are according to the indices of their vertices. The vertices are deliberately given in an unintuitive order to allow for the ordering in $\prec_{i}$ by indices and to make the (degenerate) cases behave nicely later. Let $\ell_{1},\ell_{2},\ell_{3},\ell_{4}\in[h]$ such that $\ell_{1}<\ell_{2}<\ell_{3}<\ell_{4}$. A (non-degenerate) $(\ell_{1},\ell_{2},\ell_{3},\ell_{4})$-plug, see Fig. 2, contains vertices $u_{1}$, $u_{2}$, $u_{3}$, $u_{4}$, $u_{5}$, and $u_{6}$, where $\gamma(u_{5})=\ell_{1}$, $\gamma(u_{3})=\gamma(u_{6})=\ell_{2}$, $\gamma(u_{2})=\gamma(u_{4})=\ell_{3}$, and $\gamma(u_{1})=\ell_{4}$. It contains the edges $u_{1}u_{2}$, $u_{2}u_{3}$, $u_{3}u_{4}$, $u_{4}u_{6}$, and $u_{6}u_{5}$; i.e., a plug is a path that traverses its four levels in the order $\ell_{4},\ell_{3},\ell_{2},\ell_{3},\ell_{2},\ell_{1}$. Similarly, an $(\ell_{1},\ell_{2},\ell_{3},\ell_{4})$-socket, see Fig. 2, consists of vertices $v_{1},v_{2},\dots,v_{10}$ such that $\gamma(v_{3})=\ell_{1}$, vertices $v_{2},v_{4},v_{5},v_{9}$ occupy level $\ell_{2}$, vertices $v_{1},v_{6},v_{7},v_{10}$ occupy level $\ell_{3}$, and $\gamma(v_{8})=\ell_{4}$. It contains the edges $v_{1}v_{2}$, $v_{2}v_{3}$, $v_{3}v_{5}$, $v_{5}v_{7}$, $v_{4}v_{6}$, $v_{6}v_{8}$, $v_{8}v_{10}$, and $v_{10}v_{9}$. Observe that a socket consists of two disconnected paths whose vertices interleave on levels $\ell_{2}$ and $\ell_{3}$. Let vertices $v_{1}$ and $v_{9}$ of plugs and vertices $u_{1}$ and $u_{5}$ of sockets be connecting vertices. Connecting vertices of sockets will be identified with other vertices of the construction and connecting vertices of some plugs may be connected to other plugs via additional edges.
{subfigure}
[…]
@@
Now we lift the strict inequality restriction on the levels of our gadgets. For plugs we only require $\ell_{1}\leq\ell_{2}<\ell_{3}\leq\ell_{4}$ and for sockets we require $\ell_{1}<\ell_{2}\leq\ell_{3}<\ell_{4}$ and we call a plug or a socket degenerate if it has at least one pair of repeated levels, i.e., for some $i\in[3]$, $\ell_{i}=\ell_{i+1}$. We create the degenerate plugs and sockets by contracting the edges between vertices of the repeated levels while keeping the vertex with the lower index; see Fig. 2 and 2.
Let $P$ be a plug, and let $\mathrm{com}(P)$ be the connected component of $P$ in $G$. A plug fits into a socket if the gadgets could be “weaved” as illustrated in Fig. 2. Formally, we say an $(a_{1},a_{2},a_{3},a_{4})$-plug $P$ fits into a $(s_{1},s_{2},s_{3},s_{4})$-socket $S$ when $\min_{v\in\mathrm{com}(P)}\{\gamma(v)\}\leq s_{1}$, $\max_{v\in\mathrm{com}(P)}\{\gamma(v)\}\geq s_{4}$, and $s_{1}<a_{2}<s_{2}\leq s_{3}<a_{3}<s_{4}$. In an ordered level planar drawing of $\mathcal{G}$, we say a plug $P$ links to an $(s_{1},s_{2},s_{3},s_{4})$-socket $S$ when $P$ fits into $S$, and $P$ is drawn between the connecting vertices of $S$ (that is, the edges of $P$ traversing level $s_{3}$ are to the right of $v_{1}$ and the edges of $P$ traversing level $s_{2}$ are to the left of $v_{9}$); see Fig. 2.
A defining feature of our constructions is a division into vertical strips, which is accommodated by the following notion. Given a set $L$ of levels, a wall of $L$ is a path that starts in a vertex on the bottommost level in $L$, goes through a vertex on each intermediate level of $L$, and ends at a vertex on the topmost level of $L$. Note that each gadget type essentially has a unique drawing in the sense that it corresponds to an ordered level planar graph for which all of its ordered level planar drawings have the same level planar embedding. This and further gadget properties are detailed in Appendix B. We use Lemma 3.1 to design specific plugs that can or cannot link to specific sockets in the same ordered level planar drawing.
[…]
@@
Proof 3.2 (Proof (sketch)).
To distinguish the vertices of the plugs $A$ and $B$, we use the name of the plug as superscript, so, e.g., we refer to $u_{3}^{A}$ and $u_{3}^{B}$. If the conditions are satisfied, we can draw both plugs as depicted in Fig. 3. Towards a contradiction, assuming the conditions are not satisfied, we observe that the drawing must locally “nest” plug $B$ into plug $A$ while drawing them around the edge $v_{4}v_{6}$ of $S$, at the same time $B$ “nests” into $A$ while going around the edge $v_{7}v_{5}$ of $S$; see Fig. 3. The first nesting implies that the edge $u^{B}_{3}u^{B}_{4}$ is to the left of $u^{A}_{3}u^{A}_{4}$ and the second nesting implies the opposite order – a contradiction.
{subfigure}
[…]
Figure 3: Sketches for the proof of Lemma 3.1. The edges of the socket $S$ are bold.
@@
Proof 3.4 (Proof (sketch)).
We give an overview of a parameterized reduction from MCIS to OLP; refer to Fig. 4 for an illustration. The ordered level graph constructed in this reduction admits a grid-cell structure where the vertical strips are separated by walls and the horizontal strips represent related groups of levels. More precisely, we allocate a constant number of consecutive levels per color of the MCIS instance, to which we refer as a color band. Between each pair of neighboring color bands, we have a collision band. We group several consecutive vertical strips to obtain blocks. We call the leftmost and rightmost block choice blocks and the other blocks edge blocks. Edge blocks correspond bijectively to edges of the MCIS instance.
In the grid cells of the color bands, we insert sockets and we add as many plugs as we have sockets there. No two plugs can link to the same socket, so all available sockets are occupied by one plug. Each plug is a separate connected component, which allows that plug to freely “choose” a socket within its color band where it fits. However, we have different types of plugs and sockets – the first type of plug, called high plug, fits only into the sockets of the choice blocks. In a drawing of our ordered level graph, the number of high plugs that link to the left (and not to the right) choice block within a color band forces the second type of plugs, called color plugs, to be shifted further to the right. In the MCIS instance, a shift by $i$ corresponds to selecting the $i$-th vertex of the respective color to be in the solution.
To obtain an independent set, we need to prevent that the shifts select both endpoints of an edge. To that end, we model an edge $uv$ as follows. The color plugs of one color band all occupy the same set of levels so they form a sequence specified in the ordered level graph. Because of this, we can identify, for a given shift, what socket a particular color plug links to. Consider the edge block of $uv$ and the color plugs that necessarily link to its sockets (a block has circa twice as many sockets than the maximal shift). For an illustration of the construction, assume that we have a shift in the color band of $u$ that selects $u$ and a shift in the color band of $v$ that selects $v$. We select a specific color plug of the color band of $u$ and one in the color band of $v$ to be extended so that they “reach” towards each other and end up in the same cell of a collision band. To this cell, we add a collision socket. If at most one extended plug ends up in a collision socket, then it can be drawn planarly. However, if both extended plugs end up in the same collision socket, then this cannot be drawn without crossings. Hence, there is an ordered level planar drawing of our ordered level graph if and only if there is a solution (i.e., an independent set) to the given MCIS instance.

### v2

@@ 2 An XP / XNLP Algorithm for Ordered Level Planarity
Proof.
For the “if”-direction, assume that there exists an index $1\leq j\leq h$ where $p_{j}\geq 1$ and such that $T[s_{j}^{\prime},U_{j}^{\prime}]=\texttt{true}$. Then there exists a nice sweeping sequence $S^{\prime}=(s_{1},s_{2},\dots,s_{z}=s_{j}^{\prime})$ for $\mathcal{G}$ that starts with $s_{1}=(0,0,\dots,0)$, ends with $s_{j}^{\prime}$, and uses all edges in $U_{j}^{\prime}$, as well as all edges in $E$ incident to at least one vertex to the left of $s_{j}^{\prime}$. Let $S$ denote the nice sweeping sequence obtained by appending $s$ to $S^{\prime}$. We will show that $S$ is a certificate for the fact that $T[s,U]$ is indeed true.
We start by showing that $S$ uses all edges that are incident to at least one vertex to the left of $s$. Let $e\in E$ be an edge with an endpoint $u$ to the left of $s$. If at least one endpoint of $e$ is also to the left of $s_{j}^{\prime}$, the definition of $S^{\prime}$ implies that ($S^{\prime}$ and, hence) $S$ uses $e$. So assume that no endpoint of $e$ is to the left of $s_{j}^{\prime}$. It follows that $p_{j}$ represents a gap, $p_{j-1}$ represents the vertex $u$, and the other endpoint of $e$ is on $s_{j}^{\prime}$ (it cannot lie to the right of ($s$ and, hence) $s_{j}^{\prime}$ since $T[s_{j}^{\prime},U_{j}^{\prime}]=\texttt{true}$ and, hence, $U_{j}^{\prime}\neq\bot$). Consequently, $e\in U_{j}^{\prime}$ (by definition of $U_{j}^{\prime}$) and, thus, ($S^{\prime}$ and) $S$ uses $e$, as desired.
It remains to show that $S$ uses all edges in $U$. Obviously, ($s$ and) $S$ use all edges in $U$ that join two vertices that are consecutive along $s$. So let $e\in U$ be an edge that joins two vertices that are nonconsecutive along $s$. When creating $U_{j}^{\prime}$ from $U$, only edges between consecutive vertices on $s$ are removed. Hence, $e\in U_{j}^{\prime}$ and, thus, ($S^{\prime}$ and) $S$ uses $e$, as desired.
Altogether, it follows that $S$ is indeed a certificate for the fact that $T[s,U]$ is true.
For the “only if”-direction, assume that $T[s,U]=\texttt{true}$. Then there exists a nice sweeping sequence $S=(s_{1},s_{2},\dots,s_{z}=s)$ for $\mathcal{G}$ that starts with $s_{1}=(0,0,\dots,0)$, ends with $s$, and uses all edges in $U$, as well as all edges in $E$ incident to at least one vertex to the left of $s$. Since $S$ is nice, there is an index $1\leq j\leq h$ such that
$s_{z-1}=(p_{1},p_{2},\dots,p_{j-1},p_{j}-1,p_{j+1},\dots,p_{h}),$
i.e., $s_{z-1}=s_{j}^{\prime}$. Let $S^{\prime}=(s_{1},s_{2},\dots,s_{z-1}=s^{\prime})$. We will show that $S^{\prime}$ is a certificate for the fact that $T[s_{j}^{\prime},U_{j}^{\prime}]$ is indeed true.
We begin by showing that $U_{j}^{\prime}\neq\bot$. To this end, assume otherwise, i.e., $U_{j}^{\prime}=\bot$. We distinguish two cases. First, assume that $p_{j}$ represents a vertex $v$. This implies that $U$ contains an edge $e$ joining $v$ with a vertex on $s$ that is not its predecessor or successor. The edge $e$ is not used by $s$. However, it is also not used by $S^{\prime}$ since $v$ is to the right of $s_{j}^{\prime}$; a contradiction to the fact that $S$ uses all edges in $U$. Second, assume that $p_{j}$ represents a gap and, thus, $p_{j-1}$ represents a vertex $v$. This implies that $v$ has an edge $e$ to a vertex to the right of $s$. Thus, the sequence $S$ does not use $e$. However, the endpoint $v$ of $e$ is to the left of $s$; a contradiction to the fact that $S$ uses all edges with at least one endpoint to the left of $s$. Altogether, this shows that indeed $U_{j}^{\prime}\neq\bot$.
We now show that $S^{\prime}$ uses all edges that are incident to at least one vertex to the left of $s_{j}^{\prime}$. To this end, let $u\in V$ be a vertex to the left of $s_{j}^{\prime}$. By definition of $s_{j}^{\prime}$, this vertex is also to the left of $s$. By definition of $S$, all edges incident to $u$ are used by $S$. However, none of these edges can be used by $s$ (since $u$ is to the left of $s$) and, hence, each of these edges is used by (some separation in) $S^{\prime}$, as desired.
It remains to show that $S^{\prime}$ uses all edges in $U_{j}^{\prime}$. To this end, let $e\in U_{j}^{\prime}$. First, assume that $e\in U_{j}^{\prime}\setminus U$. The definition of $U_{j}^{\prime}$ implies that $p_{j}$ represents a gap and $p_{j-1}$ represents a vertex $v$ that is incident to $e$. By definition of $S$ and the fact that $v$ is to the left of $s$, it follows that $S$ uses $e$. In fact, since $s$ cannot use $e$, it follows that (some separation in) $S^{\prime}$ uses $e$, as desired. It remains to consider the case $e\notin U_{j}^{\prime}\setminus U$. Note that this case assumption implies that both endpoints of $e$ belong to $s$. We distinguish two subcases: first assume that the endpoints of $e$ are consecutive vertices along $s$. Combining this assumption with the fact that $e\in U_{j}^{\prime}\cap U$ and the definition of $U_{j}^{\prime}$ implies that the endpoints of $e$ are also consecutive vertices along $s_{j}^{\prime}$. Hence, ($s^{\prime}$ and) $S^{\prime}$ uses $e$, as desired. For the other subcase, assume that the two endpoints of $e$ are nonconsecutive along $s$. In this case, the separation $s$ cannot use $e$ and, hence, the definition of $S$ implies that (some separation in) $S^{\prime}$ uses $e$, as desired.
Altogether, we have established that $S^{\prime}$ is indeed a certificate for the fact that $T[s_{j}^{\prime},U_{j}^{\prime}]$ is true. This concludes the proof of the claim. ∎
In view of Claim 3, we can now state our algorithms.
Deterministic construction algorithm.  Our dynamic programming table $T$ has a total of $2^{\binom{h}{2}}\prod_{j\in[h]}(2\lambda_{j}+1)$ entries. To compute the value of a table entry $T[s,U]$, we simply construct the tuples $(s_{1}^{\prime},U_{1}^{\prime}),\dots,(s_{h}^{\prime},U_{h}^{\prime})$ and then set $T[s,U]=\bigvee_{j\in[h]}T[s_{j}^{\prime},U_{j}^{\prime}]$ (the set $U_{j}^{\prime}$ is only constructed and $T[s_{j}^{\prime},U_{j}^{\prime}]$ only taken into account when $s_{j}^{\prime}$ contains no negative entry). Disregarding the time spend for the recursive calls, the time to process a table entry is clearly polynomial. Hence, by employing memoization, we can fill the table using $\mathcal{O}^{*}(2^{\binom{h}{2}}\prod_{j\in[h]}(2\lambda_{j}+1))$ time and space. To construct the solution, we can employ the usual back-linking strategy. In particular, for each true entry $T[s,U]$, it suffices to store an index $j$ such that $T[s_{j}^{\prime},U_{j}^{\prime}]$ is true, which does not change the asymptotic time/space requirements. Figure 1 illustrates a sequence of table entries corresponding to a solution.
@@
Nondeterministic decision algorithm.  We perform $2n$ steps. In each step, we nondeterministically guess the next separation $s$ and its set of prescribed edges $U$, and check whether for some $j\in[h]$, the previous separation $s^{\prime}$ is equal to $s_{j}^{\prime}$ and the previous set of prescribed edges $U^{\prime}$ is equal to $U_{j}^{\prime}$ (starting with $s^{\prime}=(0,0,\dots,0)$ and $U^{\prime}=\emptyset$). At each point in time, we only need to keep two separations and two edge sets in memory. A separation can be stored using $\mathcal{O}(h\log I)$ space. Each edge set $U$ can be encoded by means of a $h\times h$ Boolean adjacency matrix of size $\mathcal{O}(h^{2})$. The verification in each step is easy to do using polynomial time and $\mathcal{O}(\log I)$ additional space (in particular, we only need to store $\mathcal{O}(1)$ pointers / indices). ∎
As a corollary of Lemmas 1, 2 and 3, we obtain the algorithmic statements in Theorem 1:
Theorem 3.
Ordered Level Planarity parameterized by the height is XNLP. Moreover, there is a constructive XP-time algorithm for Ordered Level Planarity parameterized by the height.
@@ 3 XNLP-Hardness of Ordered Level Planarity
Our construction of $\mathcal{G}$ is heavily based on two gadgets that we call plugs and sockets (a very basic version of these gadgets was already used earlier, in an NP-hardness proof for CLP [8]; here we introduce generalized versions). We define both in terms of the list of levels their vertices occupy. Their $\prec_{i}$ orderings are according to the indices of their vertices. The vertices are deliberately given in an unintuitive order to allow for the ordering in $\prec_{i}$ by indices and to make the (degenerate) cases behave nicely later. Let $\ell_{1},\ell_{2},\ell_{3},\ell_{4}\in[h]$ such that $\ell_{1}<\ell_{2}<\ell_{3}<\ell_{4}$. A (non-degenerate) $(\ell_{1},\ell_{2},\ell_{3},\ell_{4})$-plug, see Figure 2(a), contains vertices $u_{1}$, $u_{2}$, $u_{3}$, $u_{4}$, $u_{5}$, and $u_{6}$, where $\gamma(u_{5})=\ell_{1}$, $\gamma(u_{3})=\gamma(u_{6})=\ell_{2}$, $\gamma(u_{2})=\gamma(u_{4})=\ell_{3}$, and $\gamma(u_{1})=\ell_{4}$. It contains the edges $u_{1}u_{2}$, $u_{2}u_{3}$, $u_{3}u_{4}$, $u_{4}u_{6}$, and $u_{6}u_{5}$; i.e., a plug is a path that traverses its four levels in the order $\ell_{4},\ell_{3},\ell_{2},\ell_{3},\ell_{2},\ell_{1}$. Similarly, an $(\ell_{1},\ell_{2},\ell_{3},\ell_{4})$-socket, see Figure 2(b), consists of vertices $v_{1},v_{2},\dots,v_{10}$ such that $\gamma(v_{3})=\ell_{1}$, vertices $v_{2},v_{4},v_{5},v_{9}$ occupy level $\ell_{2}$, vertices $v_{1},v_{6},v_{7},v_{10}$ occupy level $\ell_{3}$, and $\gamma(v_{8})=\ell_{4}$. It contains the edges $v_{1}v_{2}$, $v_{2}v_{3}$, $v_{3}v_{5}$, $v_{5}v_{7}$, $v_{4}v_{6}$, $v_{6}v_{8}$, $v_{8}v_{10}$, and $v_{10}v_{9}$. Observe that a socket consists of two disconnected paths whose vertices interleave on levels $\ell_{2}$ and $\ell_{3}$. Let vertices $v_{1}$ and $v_{9}$ of plugs and vertices $u_{1}$ and $u_{5}$ of sockets be connecting vertices. Connecting vertices of sockets will be identified with other vertices of the construction and connecting vertices of some plugs may be connected to other plugs via additional edges.
@@
Now we lift the strict inequality restriction on the levels of our gadgets. For plugs we only require $\ell_{1}\leq\ell_{2}<\ell_{3}\leq\ell_{4}$ and for sockets we require $\ell_{1}<\ell_{2}\leq\ell_{3}<\ell_{4}$ and we call a plug or a socket degenerate if it has at least one pair of repeated levels, i.e., for some $i\in[3]$, $\ell_{i}=\ell_{i+1}$. We create the degenerate plugs and sockets by contracting the edges between vertices of the repeated levels while keeping the vertex with the lower index; see Figures 2(c) and 2(d).
Let $P$ be a plug, and let $\mathrm{com}(P)$ be the connected component of $P$ in $G$. A plug fits into a socket if the gadgets could be “weaved” as illustrated in Figure 2(e). Formally, we say an $(a_{1},a_{2},a_{3},a_{4})$-plug $P$ fits into a $(s_{1},s_{2},s_{3},s_{4})$-socket $S$ when $\min_{v\in\mathrm{com}(P)}\{\gamma(v)\}\leq s_{1}$, $\max_{v\in\mathrm{com}(P)}\{\gamma(v)\}\geq s_{4}$, and $s_{1}<a_{2}<s_{2}\leq s_{3}<a_{3}<s_{4}$. In an ordered level planar drawing of $\mathcal{G}$, we say a plug $P$ links to an $(s_{1},s_{2},s_{3},s_{4})$-socket $S$ when $P$ fits into $S$, and $P$ is drawn between the connecting vertices of $S$ (that is, the edges of $P$ traversing level $s_{3}$ are to the right of $v_{1}$ and the edges of $P$ traversing level $s_{2}$ are to the left of $v_{9}$); see Figure 2(e).
A defining feature of our constructions is a division into vertical strips, which is accommodated by the following notion. Given a set $L$ of levels, a wall of $L$ is a path that starts in a vertex on the bottommost level in $L$, goes through a vertex on each intermediate level of $L$, and ends at a vertex on the topmost level of $L$. Note that each gadget type essentially has a unique drawing in the sense that it corresponds to an ordered level planar graph for which all of its ordered level planar drawings have the same level planar embedding. We use Lemma 4 to design specific plugs that can or cannot link to specific sockets in the same ordered level planar drawing.
Lemma 4.
@@
Proof.
To distinguish the vertices of the plugs $A$ and $B$, we use the name of the plug as superscript, so, e.g., we refer to $u_{3}^{A}$ and $u_{3}^{B}$. As we have only one socket, we use just $v_{1},\dots,v_{10}$ for the vertices of $S$. For the names of the vertices, we assume that none of the involved gadgets is degenerate. If some of the gadgets are degenerate, adjusting the vertex names is straight-forward.
(a) If the order of the layers follows the pattern $a_{2},b_{2},b_{3},a_{3}$, then there is no crossing-free drawing that has both $A$ and $B$ linked to $S$.
(b) If the order of the layers follows the pattern $a_{2},b_{2},a_{3},b_{3}$, then there is a crossing-free drawing that has both $A$ and $B$ linked to $S$.
Figure 3: Sketches for the proof of Lemma 4. The edges of the socket $S$ are bold.
For the “only if”-direction. Let an ordered level planar drawing where both $A$ and $B$ link to $S$ be given. By the definition of linking, $A$ and $B$ fit into $S$ and, hence, $a_{2},b_{2}<s_{2}\leq s_{3}<a_{3},b_{3}$. Suppose for a contradiction that $a_{2}<b_{2}$ and $a_{3}>b_{3}$; see Figure 3(a) for an illustration. (The case $a_{2}>b_{2}$ and $a_{3}<b_{3}$ is identical up to renaming $A$ and $B$.) By the definition of linking and due to the two (uncrossed) paths of $S$, the edges $u_{2}^{A}u_{3}^{A}$ and $u_{2}^{B}u_{3}^{B}$ are to the left of $v_{4}$ and $v_{6}$, while the edges $u_{3}^{A}u_{4}^{A}$ and $u_{3}^{B}u_{4}^{B}$ are to the right of $v_{4}$ and $v_{6}$. Given $a_{2}<b_{2}$, it follows that this is possible if and only if $u_{2}^{B}$, $u_{3}^{B}$, and $u_{4}^{B}$ lie (horizontally) between the edges $u_{2}^{A}u_{3}^{A}$ and $u_{3}^{A}u_{4}^{A}$. Symmetrically, the edges $u_{3}^{A}u_{4}^{A}$ and $u_{3}^{B}u_{4}^{B}$ are to the left of $v_{7}$ and $v_{5}$, while the edges $u_{4}^{A}u_{6}^{A}$ and $u_{4}^{B}u_{6}^{B}$ are to the right of $v_{7}$ and $v_{5}$. Given $a_{3}>b_{3}$, it follows that this is possible if and only if $u_{3}^{B}$, $u_{4}^{B}$, and $u_{6}^{B}$ lie between the edges $u_{3}^{A}u_{4}^{A}$ and $u_{4}^{A}u_{6}^{A}$. Therefore, $u_{3}^{B}$ and $u_{4}^{B}$ lie at the same time to the left and to the right of $u_{3}^{A}u_{4}^{A}$, which is a contradiction.
For the “if”-direction. Assume that $s_{1}<a_{2}<b_{2}<s_{2}<s_{3}<a_{3}<b_{3}<s_{4}$; see Figure 3(b) for an illustration. The case $s_{1}<b_{2}<a_{2}<s_{2}<s_{3}<b_{3}<a_{3}<s_{4}$ is identical up to renaming $A$ and $B$. We can get an ordered level planar drawing where both plugs $A$ and $B$ are to the right of $v_{1}$, and $A$ and $B$ are to the left of $v_{9}$ if we place the edges occurring between the involved layers in the following order from left to right: first $v_{1}v_{2}$ and $v_{2}v_{3}$, then $u_{2}^{A}u_{3}^{A}$, then $u_{2}^{B}u_{3}^{B}$, then $v_{4}v_{6}$ and $v_{6}v_{8}$, then $u_{3}^{B}u_{4}^{B}$, then $u_{3}^{A}u_{4}^{A}$, then $v_{3}u_{5}$ and $v_{5}v_{7}$, then $u_{4}^{A}u_{6}^{A}$, then $u_{4}^{B}u_{6}^{B}$, and finally $v_{8}u_{10}$ and $v_{10}v_{9}$. ∎
In an (ordered) level planar drawing of $\mathcal{G}$, we define spatial relations between two gadgets in the following way. For a gadget $X$, we say a level $\ell$ (or a gadget $Y$) is above $X$ if the level $\ell$ (each level hosting vertices of $Y$) has a larger index than any level hosting a vertex of $X$. Symmetrically, we define below for the levels with a smaller index. We say a gadget $Y$ is left (resp. right) of $X$ if on every level all vertices and edges of $Y$ lie to the left (resp. right) of all the vertices of $X$.
We connect plugs and sockets with other gadgets using their connecting vertices. The spatial relation of two gadgets is usually clear, so we shortcut the description of how the connection is formed in the following way. Connecting a socket $S$ to a wall $W$ is done by identifying a vertex $x$ of $S$ with the vertex of $W$ on level $\gamma(x)$. The $x=v_{1}$ if $W$ is left of $S$ and $x=v_{5}$ if $W$ is right of $S$. Connecting a plug $P$ to a gadget $Y$ means adding an edge from a connecting vertex $z$ of $P$ to $Y$. We $z=u_{1}$ if $Y$ is above $P$ and $z=u_{5}$ if $Y$ is below $P$. The other endpoint of the edge depends on $Y$. In our case, $Y$ is either a plug or a single vertex. In case $Y$ is a plug its connecting vertex is determined in the same way we did it for $P$. Hence, the connection of $P$ to $Y$ results in an edge between $u_{1}$ of $P$ and $u_{5}$ of $Y$, or vice-versa. See connecting vertices of plugs and sockets in Figure 2.
An ordered level planar graph is called rigid if all of its ordered level planar drawings have a common level planar embedding. Recall that $\mathcal{G}$ is proper if $\gamma(v)=\gamma(u)+1$ for every edge $(u,v)\in E(G)$.
Observation 1.
Every proper ordered level planar graph is rigid.
Observation 2.
A $(\ell_{1},\ell_{2},\ell_{3},\ell_{4})$-plug or a $(\ell_{1},\ell_{2},\ell_{3},\ell_{4})$-socket correspond to a proper ordered level graph if and only if $\ell_{i+1}\in\{\ell_{i},\ell_{i}+1\}$ for every $i\in[3]$.
By inserting a level $i$ into an ordered level graph $\mathcal{G}$, we mean that the number $h$ of levels of $\mathcal{G}$ increases by one and, for each vertex $v$ of $\mathcal{G}$ with $\gamma(v)\geq i$, its level $\gamma(v)$ increases by one.
Observation 3.
A rigid ordered level graph remains rigid even if we insert new levels. Moreover, a rigid graph $\mathcal{G}$ remains rigid even if we subdivide an edge $uv$ by a new vertex $w$, i.e., if we replace the directed edge $uv$ by the directed edges $uw$ and $wv$, and set the x-coordinate of $w$ according to where $uv$ crossed $L_{\gamma(w)}$ in the unique drawing of $\mathcal{G}$.
From Observations 1, 2, and 3, it follows that plugs, sockets, and walls are all rigid since we can start with a proper ordered level graph and then insert additional levels.
We are now done with the auxiliary definitions. We first give the full proof of $W[1]$-hardness of OLP. Then we extend the construction to show XNLP-hardness of OLP. Last in this section, we show that the XNLP-hardness result also works for connected ordered level graphs.
@@
Proof.
Let $\mathcal{S}$ be an instance of MCIS, that is, a graph $H$, an integer $k$, and a $k$-partition $C_{1},C_{2},\dots,C_{k}$ of $V(H)$. We prove the theorem by constructing an OLP instance $\mathcal{G}$, with $h\in\mathcal{O}(k)$ levels, that has an ordered level planar drawing if and only if $\mathcal{S}$ contains a multicolored independent set of size $k$.
Let $n=|V(H)|$ and $m=|E(H)|$. For a color $j\in[k]$, let $m_{j}=|\{uv\colon c(u)=j\lor c(v)=j\}|$, i.e., $m_{j}$ is the number of edges incident to a vertex with color $j$. Note that $\sum_{j=1}^{k}m_{j}=2m$ by double counting since we assume that the colors of the endpoints of every edge are distinct. For this proof, we also assume that, without loss of generality, the color classes have the same size $n^{\prime}$, i.e., $n^{\prime}=|C_{j}|=n/k$ for each $j\in[k]$.